# Adaptive Prompt: Unlocking the Power of Visual Prompt Tuning

### v4

@@ 6 Experiment
In this section, we compare VAPT with VPT and other widely used PEFT methods. We also examine the robustness of VAPT under different pre-training objectives and present our findings on sample efficiency. For additional results, including ablation studies, computational cost, and interpretative visualizations, please refer to Appendix E.
@@ 6.1 Experimental Setup
Datasets. We evaluate our approach on two benchmarks: FGVC and VTAB-1K [75]. FGVC includes five fine-grained classification datasets: CUB [66], Oxford Flowers [50], Stanford Cars [14], Stanford Dogs [31], and NABirds [63], which challenge models to distinguish visually similar classes. VTAB-1K comprises 19 datasets, each with 1,000 training examples, grouped into three categories: Natural (images captured by standard cameras), Specialized (images collected via specialized equipment), and Structured (tasks requiring structural understanding, such as 3D depth prediction).
Baselines. To ensure consistency with prior work [28, 17], we compare VAPT against commonly used PEFT methods. Our experiments include standard Vision Transformer (ViT) [8] architectures with supervised pre-trained on ImageNet-21K [7]. Additionally, we evaluate the performance of VAPT on two self-supervised learning objectives: MAE [20] and MoCo v3 [4].
Training. We perform a grid search on val set of each task to determine the optimal learning rate, weight decay, kernel size $K$, and projector dimension $r$. We schedule the learning rate with cosine decay and train for 100 epochs. Following [28], we use batch sizes of 64 and 128. Additional details can be found in Appendix D.
@@ 6.2 Empirical Results
Pre-trained objectives MAE [20] MoCo v3 [4]
Tuned/ VTAB-1K [75] [19] Tuned/ VTAB-1K [75] [19]
@@
Full [26] 100.00 59.31 79.68 53.82 100.00 71.95 84.72 51.98
Linear [26] 0.04 18.87 [0] 53.72 [0] 23.70 [0] 0.04 67.46 [4] 81.08 [0] 30.33 [0]
Partial-1 [74] 8.30 58.44 [5] 78.28 [1] 47.64 [1] 8.30 72.31 [5] 84.58 [2] 47.89 [1]
Bias [56] 0.16 54.55 [1] 75.68 [1] 47.70 [0] 0.16 72.89 [3] 81.14 [0] 53.43 [4]
Adapter [2] 0.87 54.90 [3] 75.19 [1] 38.98 [0] 1.12 74.19 [4] 82.66 [1] 47.69 [2]
VPT-Shallow [28] 0.05 39.96 [1] 69.65 [0] 27.50 [0] 0.06 67.34 [3] 82.26 [0] 37.55 [0]
VPT-Deep [28] 0.31 36.02 [0] 60.61 [1] 26.57 [0] 0.22 70.27 [4] 83.04 [0] 42.38 [0]
GateVPT [73] 0.05 47.61 [2] 76.86 [1] 36.80 [1] 0.06 74.84 [4] 83.38 [1] 49.10 [3]
@@
Overall comparison. In Table 1, we present the average accuracy on VTAB-1K and FGVC, comparing VAPT with full fine-tuning and other widely used PEFT methods. Specifically, Full [26] denotes full fine-tuning, which updates both backbone and classification head, whereas Linear [26], Partial-1 [74] (top layer), and MLP-3 [3] (3 MLP layers) only update partial parameters. Sidetune [76], Bias [56], Adapter [2], and LoRA [25] introduce trainable modules for adaptation. VPT [28] and E2VPT [17] are concurrent visual prompt tuning approaches (see Appendix C). We observe that VAPT outperforms full fine-tuning on across 22 out of 24 tasks. Specifically, it achieves a notable 1.04% increase in accuracy on FGVC and an substantial 11.70% improvement on VTAB-1K Structured. On VTAB-1K, VAPT yields a 7.34% average gain over full fine-tuning, while updating only 0.36% of the backbone parameters. These results underscore both the effectiveness and efficiency of VAPT as an innovative PEFT method. Additionally, VAPT achieves state-of-the-art performance compared to other PEFT approaches. Among prompt tuning methods, VAPT consistently demonstrates its superiority, surpassing VPT in 23 out of 24 tasks despite utilizing fewer parameters. We attribute these improvements to VAPT’s simple but elegant design, which leverages input information to enhance the expressiveness of prompt experts. These findings highlight the potential of adaptive prompts to improve performance while reducing parameter usage.
@@
Different pre-training methods. In our experiments, we investigate the impact of self-supervised pre-training methods by evaluating two objectives: MAE [20] and MoCo v3 [4], as detailed in Table 2. Previous studies [28, 73] have highlighted the relatively suboptimal performance of VPT when applied to backbones pre-trained using self-supervised objectives. In contrast, VAPT achieves significant performance improvements by effectively leveraging the rich information present in the input features. For example, VAPT demonstrates a remarkable 23.21% improvement in accuracy under MAE on VTAB-1K Natural tasks and an 18.36% improvement in accuracy under MoCo v3 on VTAB-1K Structured. Furthermore, compared to other PEFT methods, VAPT consistently outperforms them, achieving the highest “Number of Wins” relative to full fine-tuning, with 9 out of 19 instances under MAE and 15 out of 19 instances under MoCo v3. These results underscore the generality and robustness of our method, both theoretically (see Section 5) and empirically across different pre-training objectives.
Figure 4: VAPT outperforms the VPT in sample efficiency.
Sample efficiency. As discussed in Section 5, our formulation achieves optimal sample efficiency for prompt estimation. To validate these theoretical insights, we systematic experiments on the Stanford Dogs dataset [31]. Following [10], we subsample each class by fraction $f=\{0.01,0.1,0.3,0.5,1.0\}$ and scaling the number of training epochs by $1/f$ so that the total number of images presented to the model remains constant. The results in Figure 4 and Table 3 show that VAPT consistently outperforms VPT across all training set sizes. Notably, with only 1% of the data, VPT obtains a modest accuracy of 3.6%, whereas VAPT achieves a substantially higher accuracy of 60.1%. Furthermore, to match the performance of VPT trained on the full dataset, VAPT requires only 30% of the data, yielding an approximate 3×3$\times$3 × reduction. This highlights the superior sample efficiency of our approach.
Table 3: Classification accuracy on the Stanford Dogs [31] dataset, where a variable fraction of images per class is retained during training. Bold highlights the best results.
@@ 7 Conclusion
In this paper, we highlight the restricted functional expressiveness in existing VPT prompt formulations and introduce VAPT, a novel adaptive prompt design. Our approach demonstrates superior performance compared to VPT while using fewer parameters, and achieves optimal sample efficiency both theoretically and empirically. However, our current investigation is limited to implementations on the Vision Transformer architecture and evaluations on classification tasks. Given the promising potential of adaptive prompts, future research could explore alternative designs across different transformer architectures and extend their application to a broader range of tasks. Additionally, future work may investigate the potential synergies between VAPT and VPT through their integration.
Supplement to “Adaptive Prompt: Unlocking the Power of
Visual Prompt Tuning”
In this supplementary material, we provide detailed proofs of the main results in Appendix A, along with additional theoretical findings in Appendix B. A discussion of related work can be found in Appendix C. Implementation details of our experiments are described in Appendix D, while additional experimental results are included in Appendix E. Specifically, per task detail results corresponding to the main experiments in Section 6 are reported in Appendix E.1 and Appendix E.2. The performance of VAPT across different backbone scales is evaluated in Appendix E.3. Ablation studies, computational cost analyses, and interpretive visualizations are presented in Appendix E.4, Appendix E.5, and Appendix E.6, respectively.
@@ Appendix A Proofs
In this appendix, we provide proofs for key theoretical results for prompt estimation in the main text.
@@ A.1 Proof of Theorem 1
Given the least square estimator $\widehat{G}_{n}$ in equation (18), under the $L_{2}(\mu)$ norm, the estimator $f_{\widehat{G}_{n}}(\cdot)$ converges to the true model $f_{G_{*}}(\cdot)$ at a parametric rate in the sample size. That is,
@@
We divide the proof of the above inequality into local and global parts respectively in Appendices A.1.1 and A.1.2. Before going into the proof details, let us introduce some essential assumptions on the activation function $\sigma$.
@@
(A.1) (Identifiability) If there exist parameters $(W_{1},W_{2})$ and $(W_{1}^{\prime},W_{2}^{\prime})$ such that $W_{2}\sigma(W_{1}{\bm{X}})=W_{2}^{\prime}\sigma(W_{1}^{\prime}{\bm{X}})$ for almost surely ${\bm{X}}$, then we obtain that $(W_{1},W_{2})=(W_{1}^{\prime},W_{2}^{\prime})$.
@@
for any vector $\gamma\in\mathbb{R}^{2dr}$ and for some positive constants $\zeta$ and $C$ which are independent of ${\bm{X}}$ and $(W_{1},W_{2}),(W_{1}^{\prime},W_{2}^{\prime})$. Here, $\alpha=(\alpha_{1},\alpha_{2})\in\mathbb{N}^{r\times d}\times\mathbb{N}^{d\times r}$.
(A.3) (Strong identifiability) The function $\sigma$ is twice differentiable almost surely and for any natural number $\ell$ and distinct parameters $\{W_{1,j}:j\in[\ell]\}$, the functions in the set
@@
are linearly independent for almost surely ${\bm{X}}$, where the notations $\sigma^{(1)}$ and $\sigma^{(2)}$ stand for the first and second derivatives of the function $\sigma$, which are applied element-wise to the matrices $W_{1,j}{\bm{X}}$.
A.1.1 Local part.
@@
We assume that the above inequality does not hold. Then, we can find a sequence of measures $G_{n}:=\sum_{j^{\prime}=1}^{L^{\prime}}\exp(b_{n,j^{\prime}})\delta_{(W_{n,1j^{\prime}},W_{n,2})}$ in $\mathcal{G}_{L^{\prime}}(\Theta)$ such that
@@
From the hypothesis, since $\mathcal{D}_{1n}\to 0$ as $n\to\infty$, we have $\sum_{i\in\mathcal{V}_{j}}\exp(b_{n,i})\to\exp(b_{*,j})$, $W_{n,1i}\to W_{*,1j^{\prime}}$, and $W_{n,2}\to W_{*,2}$ for any $i\in\mathcal{V}_{j},j\in[L]$. Throughout this proof, for the simplicity of the argument we assume without loss of generality that $B=I_{d}$, $C=I_{d}$, and $r=1$ with a note that our techniques can be generalized to the general case of these given matrices. Now, we divide the proof of the local part into three main substeps as follows:
@@
To streamline the argument, we define the following functions: $\bar{E}({\bm{X}};W_{1},W_{2}):=\exp((W_{2}\sigma(W_{1}{\bm{X}}))^{\top}{\bm{X}})$ and $\bar{H}({\bm{X}};W_{1},W_{2})=W_{2}\sigma(W_{1}{\bm{X}})$. Furthemore, the product of these functions is defined as $\bar{F}({\bm{X}};W_{1},W_{2})=\bar{E}({\bm{X}};W_{1},W_{2})\bar{H}({\bm{X}};W_{1},W_{2})$. To account for the difference in the number of elements among Voronoi cells, we further decompose the function $A_{n}({\bm{X}})$ as follows:
@@
An application of the first-order Taylor expansion leads to
@@
Here, the indices $i,j$ in these equations satisfy that $|\mathcal{V}_{j}|=1$, i.e., Voronoi cells with exactly one element and $i\in\mathcal{V}_{j}$. Furthermore, the functions $\bar{R}_{ij,1}({\bm{X}})$ and $\bar{R}_{ij,2}({\bm{X}})$ in these equations correspond to Taylor remainders when expanding the functions $\bar{E}$ and $\bar{H}$. Combining the above results leads to
@@
[…]
@@
We now move to decompose the function $A_{n,2}({\bm{X}})$. Different from the function $A_{n,1}({\bm{X}})$ where we utilize only first-order Taylor expansion, to account for more than one element in the Voronoi cells in $A_{n,2}({\bm{X}})$, we resort to the second-order Taylor expansions to the functions $\bar{E}$ and $\bar{H}$. In particular, an application of the second-order Taylor expansion leads to
@@
where $\alpha=(\alpha_{1},\alpha_{2})$, $\beta=(\beta_{1},\beta_{2})$. Furthermore, due to the uniform smoothest of the functions $\bar{E}$ and $\bar{H}$, the function $\tilde{R}_{n,2}({\bm{X}})$, which is the combination of Taylor remainders from the second-order Taylor expansion, satisfies that $\bar{R}_{n,2}({\bm{X}})/\mathcal{D}_{1n}\to 0$ when $n$ goes to infinity. Furthermore, the coefficients $\bar{M}_{n,j,\alpha_{1},\alpha_{2}}$ and $\bar{M}_{n,j,\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}}$ in the above formulation takes the following forms:
@@
for any index $\alpha$ such that $|\alpha|=2$ and
@@
for any coefficients $\alpha$ and $\beta$ such that $|\alpha|=|\beta|=1$. Given the formulations of the functions $\bar{E}({\bm{X}};W_{1},W_{2})$ and $\bar{H}({\bm{X}};W_{1},W_{2})$, their partial derivatives take the following forms:
@@
Similar to our decomposition of the function $A_{n}({\bm{X}})$, we also separate the Voronoi cells with exactly one element and with more than one element in the decomposition of the function $B_{n}({\bm{X}})$. Therefore, we obtain that
@@
For Voronoi cells with exactly one element, we will utilize Taylor expansion up to the first order while for Voronoi cells with more than one element, the second-order Taylor expansion will be employed. This strategy leads to the following representations:
@@
In these expressions, the functions $\bar{R}_{n,3}({\bm{X}})$ and $\bar{R}_{n,4}({\bm{X}})$ correspond to the Taylor remainders. Due to the uniform smoothness of the functions $\bar{E}$, we obtain that $R_{n,3}({\bm{X}})/\mathcal{D}_{1n}\to 0$ and $R_{n,4}({\bm{X}})/\mathcal{D}_{1n}\to 0$ as $n$ approaches infinity. By computing the closed-form formulations for the partial derivatives of the function $\bar{E}$, both the functions $B_{n,1}({\bm{X}})$ and $B_{n,2}({\bm{X}})$ can be rewritten as follows:
@@
We proceed to demonstrate that not all of the coefficients of these linear independent functions go to 0 as $n$ goes to infinity. We prove that claim by contradiction. We assume that all these coefficients go to 0 as $n$ goes to $\infty$. From the representation of the ratio $Q_{n}({\bm{X}})/\mathcal{D}_{1n}$ in terms of these linear independent functions, we obtain that these ratios $L_{n,1,j}^{(u)}/\mathcal{D}_{1n}$, $L_{n,2,j}^{(u)}/\mathcal{D}_{1n}$, $L_{n,3,j}^{(u)}/\mathcal{D}_{1n}$, $L_{n,4,j}^{(uv)}/\mathcal{D}_{1n}$, $L_{n,5,j}^{(uv)}/\mathcal{D}_{1n}$, $L_{n,6,j}^{(uv)}/\mathcal{D}_{1n}$, and $M_{n,j,\mathbf{0}_{d},\mathbf{0}_{d}}/\mathcal{D}_{1n}$ go to 0 as $n$ approaches infinity for any indices $u,v\in[d]$ and $j\in[L]$.
By first considering the vanishing of the ratio $\bar{M}_{n,j,\mathbf{0}_{d},\mathbf{0}_{d}}/\mathcal{D}_{1n}$ to 0 for any $j\in[L]$ and takes the absolute value of that ratio, we find that
@@
By varying the index $j$ from 1 to $L$ in these limits and summing these limits, we find that
@@
Our strategy now is to consider the limits corresponding to Vornoi cells with exactly one element and more than one element separately. In particular, for Voronoi cells with exactly one element, namely, for indices $j\in[L]$ such that their corresponding Voronoi cells have one element, i.e., $|\mathcal{V}_{j}|=1$, as the ratio $L_{n,3,j}^{(u)}/\mathcal{D}_{1n}$ approaches 0, we have
@@
Similarly, $L_{n,2,j}/\mathcal{D}_{1n}\to 0$ also leads to $\dfrac{\sum_{i\in\mathcal{V}_{j}}\exp(b_{n,i})\|\Delta W_{n,2ij}\|_{1}}{\mathcal{D}_{1n}}\to 0$. Combining these results leads to

### v5

@@ 6 Experiment
In this section, we compare VAPT with VPT and other widely used PEFT methods. We also examine the robustness of VAPT under different pre-training objectives and present our findings on its sample efficiency. For additional results, including ablation studies, semantic segmentation results, statistical significance tests, computational cost, and interpretive visualizations, please refer to Appendix E.
@@ 6.1 Experimental Setup
Datasets. We evaluate our approach on two benchmarks: FGVC and VTAB-1K zhai2019large . FGVC includes five fine-grained classification datasets: CUB wah2011caltech , Oxford Flowers nilsback2008automated , Stanford Cars gebru2017fine , Stanford Dogs khosla2011novel , and NABirds van2015building , which challenge models to distinguish visually similar classes. VTAB-1K comprises 19 datasets, each with 1,000 training examples, grouped into three categories: Natural (images captured by standard cameras), Specialized (images collected via specialized equipment), and Structured (tasks requiring structural understanding, such as 3D depth prediction). Beyond classification, we also assess performance on semantic segmentation using ADE20K zhou2019semantic .
Baselines. In line with prior work jia2022visual ; han20232vpt , we compare VAPT against commonly used PEFT methods. All classification experiments use standard Vision Transformer (ViT) dosovitskiy2020image that are supervised pre-trained on ImageNet-21K deng2009imagenet . For semantic segmentation, we employ SETR-PUP zheng2021rethinking , which uses ViT as the backbone encoder. Additionally, we evaluate VAPT with backbones pre-trained using two self-supervised learning: MAE he2022masked and MoCo v3 chen2021empirical .
Training. We perform a grid search on val set of each task to determine the optimal learning rate, weight decay, kernel size $K$, and projector dimension $r$. We schedule the learning rate using a cosine decay schedule and train all models for 100 epochs. Following jia2022visual , we use batch sizes of 64 and 128. Additional implementation details can be found in Appendix D.
@@ 6.2 Empirical Results
Pre-trained objectives MAE he2022masked MoCo v3 chen2021empirical
Tuned/ VTAB-1K zhai2019large [19] Tuned/ VTAB-1K zhai2019large [19]
@@
Full iofinova2022well 100.00 59.31 79.68 53.82 100.00 71.95 84.72 51.98
Linear iofinova2022well 0.04 18.87 [0] 53.72 [0] 23.70 [0] 0.04 67.46 [4] 81.08 [0] 30.33 [0]
Partial-1 yosinski2014transferable 8.30 58.44 [5] 78.28 [1] 47.64 [1] 8.30 72.31 [5] 84.58 [2] 47.89 [1]
Bias rebuffi2017learning 0.16 54.55 [1] 75.68 [1] 47.70 [0] 0.16 72.89 [3] 81.14 [0] 53.43 [4]
Adapter cai2020tinytl 0.87 54.90 [3] 75.19 [1] 38.98 [0] 1.12 74.19 [4] 82.66 [1] 47.69 [2]
VPT-Shallow jia2022visual 0.05 39.96 [1] 69.65 [0] 27.50 [0] 0.06 67.34 [3] 82.26 [0] 37.55 [0]
VPT-Deep jia2022visual 0.31 36.02 [0] 60.61 [1] 26.57 [0] 0.22 70.27 [4] 83.04 [0] 42.38 [0]
GateVPT yoo2023improving 0.05 47.61 [2] 76.86 [1] 36.80 [1] 0.06 74.84 [4] 83.38 [1] 49.10 [3]
@@
Overall Comparison. Table 1 compares VAPT’s performance against full fine-tuning and other prominent PEFT methods on VTAB-1K and FGVC. Full iofinova2022well denotes full fine-tuning, which updates all model parameters, while methods such as Linear iofinova2022well , Partial-1 yosinski2014transferable (top layer), and MLP-3 chen2020improved (3 MLP layers) modify only a subset of parameters. Sidetune zhang2020side , Bias rebuffi2017learning , Adapter cai2020tinytl , and LoRA hu2021lora introduce trainable modules for adaptation. Concurrent visual prompt tuning approaches, VPT jia2022visual and E2VPT han20232vpt , are also included for comparison (see Appendix C for further details). Our results indicate that VAPT surpasses full fine-tuning in 22 out of 24 tasks. Specifically, VAPT achieves a notable 1.04% accuracy increase on FGVC and a substantial 11.70% improvement on VTAB-1K Structured. Across the entire VTAB-1K benchmark, VAPT demonstrates an average gain of 7.34% over full fine-tuning, while updating merely 0.36% of the backbone parameters. These findings highlight VAPT’s significant effectiveness and efficiency as an innovative PEFT method. Additionally, VAPT achieves state-of-the-art performance compared to other PEFT approaches. Among prompt tuning methods, VAPT consistently outperforms VPT in 23 out of 24 tasks, despite utilizing fewer parameters. We attribute these improvements to VAPT’s effective design, which enhances the expressiveness of prompt experts. These results underscore VAPT’s potential as a powerful tool for improving model performance with significantly reduced parameter overhead.
@@
Different Pre-training Methods. We investigate VAPT’s performance when initialized with backbones pre-trained using different self-supervised learning (SSL) objectives, specifically MAE he2022masked and MoCo v3 chen2021empirical , with detailed results in Table 2. Previous studies jia2022visual ; yoo2023improving have indicated that VPT can exhibit suboptimal performance with SSL pre-trained backbones. In contrast, VAPT achieves significant performance improvements by effectively leveraging the rich information present in the input features. For example, VAPT demonstrates a remarkable 23.21% accuracy improvement under MAE on VTAB-1K Natural tasks and an 18.36% improvement with MoCo v3 on VTAB-1K Structured. Furthermore, when compared to other PEFT methods, VAPT consistently outperforms them, achieving the highest “Number of Wins” relative to full fine-tuning, with 9 out of 19 tasks under MAE and 15 out of 19 tasks under MoCo v3. These results underscore the generality and robustness of our method across different pre-training objectives, supported by both theoretical analysis (Section 5) and empirical evidence.
Figure 2: Comparison of VAPT and VPT across varying fractions of training data on Stanford Dogs.
Table 3: Classification accuracy on Stanford Dogs khosla2011novel using varying fractions of training data. Bold indicates the best performance.
@@
Sample Efficiency. As theoretically established in Section 5, our proposed formulation exhibits optimal sample efficiency for prompt estimation. To empirically validate this, we systematic experiments on the Stanford Dogs dataset khosla2011novel . Following d2021convit , we subsampled each class by fraction $f=\{0.01,0.1,0.3,0.5,1.0\}$ and scaled the number of training epochs by $1/f$, thereby ensuring a constant total number of image presentations to the model. The results, presented in Figure 2 and Table 3, demonstrate that VAPT consistently outperforms VPT across all evaluated training set sizes. Notably, with only 1% of the data, VPT achieves an accuracy of 3.6%, whereas VAPT attains a substantially higher accuracy of 60.1%. Furthermore, VAPT requires only 30% of the data to match the performance of VPT trained on the full 100% dataset, signifying an approximate 3×3$\times$3 × reduction in data requirements. This underscores the superior sample efficiency of our approach.
@@ 7 Conclusion
In this paper, we highlight the limited functional expressiveness of prompt experts in existing VPT formulation and introduce VAPT, a novel adaptive prompt design. Our approach achieves superior performance compared to VPT while using fewer parameters and demonstrates optimal sample efficiency both theoretically and empirically. Our theoretical analysis is grounded in interpreting self-attention as a composition of multiple MoE models. Given that self-attention is central to most Transformer architectures, we believe our analysis can naturally extend to other Transformer variants. While our experiments were conducted on Vision Transformers, the potential demonstrated by adaptive prompts suggests that future work could explore alternative designs across diverse Transformer architectures and broaden their applications to a wider range of tasks. Furthermore, future research could investigate the potential synergies between VAPT and VPT, for instance, by exploring their integration.
Supplement to “On the Expressiveness of Visual Prompt Experts”
In this supplementary material, we provide detailed proofs of the main results in Appendix A and additional theoretical findings in Appendix B. A discussion of related work is presented in Appendix C. Implementation details of our experiments are described in Appendix D, while additional experimental results are included in Appendix E. Specifically, these include detailed per-task results corresponding to the main experiments in Section 6 (Appendices E.1, E.2), statistical significance tests (Appendix E.3), an evaluation of VAPT performance across different backbone scales (Appendix E.4), and semantic segmentation results (Appendix E.5). Finally, ablation studies, computational cost analyses, and interpretive visualizations are presented in Appendix E.6, Appendix E.7, and Appendix E.8, respectively.
@@ Appendix A Proofs
In this appendix, we provide proofs for key theoretical results on prompt estimation presented in the main text.
@@ A.1 Proof of Theorem 1
Given the least-squares estimator $\widehat{G}_{n}$ defined in Equation (18), under the $L_{2}(\mu)$ norm, the estimator $f_{\widehat{G}_{n}}(\cdot)$ converges to the true model $f_{G_{*}}(\cdot)$ at a parametric rate with respect to the sample size. That is,
@@
We divide the proof of the above inequality into local and global parts presented in Appendices A.1.1 and A.1.2, respectively. Before presenting the proof details, let us introduce some essential assumptions on the activation function $\sigma$.
@@
(A.1) (Identifiability) If there exist parameters $(W_{1},W_{2})$ and $(W_{1}^{\prime},W_{2}^{\prime})$ such that $W_{2}\sigma(W_{1}{\bm{X}})=W_{2}^{\prime}\sigma(W_{1}^{\prime}{\bm{X}})$ holds for almost surely ${\bm{X}}$, then $(W_{1},W_{2})=(W_{1}^{\prime},W_{2}^{\prime})$.
@@
for any vector $\gamma\in\mathbb{R}^{2dr}$ and for some positive constants $\zeta$ and $C$ that are independent of ${\bm{X}}$ and $(W_{1},W_{2}),(W_{1}^{\prime},W_{2}^{\prime})$. Here, $\alpha=(\alpha_{1},\alpha_{2})\in\mathbb{N}^{r\times d}\times\mathbb{N}^{d\times r}$.
(A.3) (Strong identifiability) The function $\sigma$ is twice differentiable almost surely. For any natural number $\ell$ and distinct parameters $\{W_{1,j}:j\in[\ell]\}$, the functions in the set
@@
are linearly independent for almost surely ${\bm{X}}$. Here, $\sigma^{(1)}$ and $\sigma^{(2)}$ denote the first and second derivatives of $\sigma$, respectively, which are applied element-wise to the matrices $W_{1,j}{\bm{X}}$.
A.1.1 Local Part
@@
We assume, for the sake of contradiction, that the above inequality does not hold. Then, we can find a sequence of measures $G_{n}:=\sum_{j^{\prime}=1}^{L^{\prime}}\exp(b_{n,j^{\prime}})\delta_{(W_{n,1j^{\prime}},W_{n,2})}$ in $\mathcal{G}_{L^{\prime}}(\Theta)$ such that
@@
From the hypothesis, since $\mathcal{D}_{1n}\to 0$ as $n\to\infty$, we have $\sum_{i\in\mathcal{V}_{j}}\exp(b_{n,i})\to\exp(b_{*,j})$, $W_{n,1i}\to W_{*,1j^{\prime}}$, and $W_{n,2}\to W_{*,2}$ for any $i\in\mathcal{V}_{j},j\in[L]$. Throughout this proof, for simplicity of argument we assume without loss of generality that $B=I_{d}$, $C=I_{d}$, and $r=1$. We note that our techniques can be generalized to the general case of these given matrices. Now, we divide the proof of the local part into the following three main substeps:
@@
To streamline the argument, we define the following functions: $\bar{E}({\bm{X}};W_{1},W_{2}):=\exp((W_{2}\sigma(W_{1}{\bm{X}}))^{\top}{\bm{X}})$ and $\bar{H}({\bm{X}};W_{1},W_{2})=W_{2}\sigma(W_{1}{\bm{X}})$. Furthermore, the product of these functions is defined as $\bar{F}({\bm{X}};W_{1},W_{2})=\bar{E}({\bm{X}};W_{1},W_{2})\bar{H}({\bm{X}};W_{1},W_{2})$. To account for the difference in the number of elements among Voronoi cells, we further decompose the function $A_{n}({\bm{X}})$ as follows:
@@
An application of the first-order Taylor expansion leads to:
@@
Here, the indices $i,j$ in the above equations satisfy $|\mathcal{V}_{j}|=1$, i.e., Voronoi cells with exactly one element and $i\in\mathcal{V}_{j}$. Furthermore, the functions $\bar{R}_{ij,1}({\bm{X}})$ and $\bar{R}_{ij,2}({\bm{X}})$ in these expressions represent the Taylor remainders from the expansions of the functions $\bar{E}$ and $\bar{H}$. Combining these results leads to:
@@
where $\alpha=(\alpha_{1},\alpha_{2})$. Furthermore, due to the uniform smoothness of the functions $\bar{E}$ and $\bar{H}$, the function $\tilde{R}_{n,1}({\bm{X}})$ in the above equation satisfies $\bar{R}_{n,1}({\bm{X}})/\mathcal{D}_{1n}\to 0$ when $n$ approaches infinity. Finally, the terms $M_{n,j,\alpha_{1},\alpha_{2}}$ in this equation admit the following forms:
@@
We now proceed to decompose the function $A_{n,2}({\bm{X}})$. Unlike the function $A_{n,1}({\bm{X}})$ for which we utilized only a first-order Taylor expansion, the analysis of $A_{n,2}({\bm{X}})$ must account for Voronoi cells potentially containing more than one element. Consequently, we employ second-order Taylor expansions of the functions $\bar{E}$ and $\bar{H}$. In particular, an application of the second-order Taylor expansion leads to:
@@
where $\alpha=(\alpha_{1},\alpha_{2})$, $\beta=(\beta_{1},\beta_{2})$. Furthermore, due to the uniform smoothness of the functions $\bar{E}$ and $\bar{H}$, the function $\tilde{R}_{n,2}({\bm{X}})$, which represents the combination of Taylor remainders from the second-order Taylor expansion, satisfies $\bar{R}_{n,2}({\bm{X}})/\mathcal{D}_{1n}\to 0$ as $n$ goes to infinity. Additionally, the coefficients $\bar{M}_{n,j,\alpha_{1},\alpha_{2}}$ and $\bar{M}_{n,j,\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}}$ in the above formulation take the following forms:
@@
for any multi-index $\alpha$ such that $|\alpha|=2$ and
@@
for any coefficients $\alpha$ and $\beta$ such that $|\alpha|=|\beta|=1$. Given the expressions for the functions $\bar{E}({\bm{X}};W_{1},W_{2})$ and $\bar{H}({\bm{X}};W_{1},W_{2})$, their partial derivatives take the following forms:
@@
Similar to our decomposition of the function $A_{n}({\bm{X}})$, in decomposing the function $B_{n}({\bm{X}})$, we also distinguish between Voronoi cells containing exactly one element and those containing more than one element. Therefore, we obtain:
@@
For Voronoi cells with exactly one element, we utilize the first-order Taylor expansion, while for those with more than one element, the second-order Taylor expansion is employed. This strategy leads to the following representations:
@@
In these expressions, the functions $\bar{R}_{n,3}({\bm{X}})$ and $\bar{R}_{n,4}({\bm{X}})$ correspond to the Taylor remainders. Due to the uniform smoothness of the functions $\bar{E}$, we obtain $R_{n,3}({\bm{X}})/\mathcal{D}_{1n}\to 0$ and $R_{n,4}({\bm{X}})/\mathcal{D}_{1n}\to 0$ as $n$ approaches infinity. By computing the closed-form expressions for the partial derivatives of the function $\bar{E}$, both functions $B_{n,1}({\bm{X}})$ and $B_{n,2}({\bm{X}})$ can be rewritten as follows:
@@
Combining these results, the function $Q_{n}({\bm{X}})$ can be rewritten as follows:
@@
We now demonstrate that not all of the coefficients of these linearly independent functions converge to 0 as $n$ goes to infinity. To prove this by contradiction, assume that all these coefficients do converge to 0 as $n$ goes to $\infty$. From the representation of the ratio $Q_{n}({\bm{X}})/\mathcal{D}_{1n}$ in terms of these linearly independent functions, it follows that these ratios $L_{n,1,j}^{(u)}/\mathcal{D}_{1n}$, $L_{n,2,j}^{(u)}/\mathcal{D}_{1n}$, $L_{n,3,j}^{(u)}/\mathcal{D}_{1n}$, $L_{n,4,j}^{(uv)}/\mathcal{D}_{1n}$, $L_{n,5,j}^{(uv)}/\mathcal{D}_{1n}$, $L_{n,6,j}^{(uv)}/\mathcal{D}_{1n}$, and $M_{n,j,\mathbf{0}_{d},\mathbf{0}_{d}}/\mathcal{D}_{1n}$ all converge to 0 as $n$ approaches infinity for all indices $u,v\in[d]$ and $j\in[L]$.
By first considering the vanishing of the ratio $\bar{M}_{n,j,\mathbf{0}_{d},\mathbf{0}_{d}}/\mathcal{D}_{1n}$ to 0 for any $j\in[L]$ and then taking the absolute value of this ratio, we find that
@@
By varying the index $j$ from 1 to $L$ and summing the corresponding limits, we find that
@@
Our strategy is now to consider separately the terms corresponding to Voronoi cells with exactly one element and those with more than one element. In particular, for Voronoi cells with exactly one element, i.e., for indices $j\in[L]$ such that $|\mathcal{V}_{j}|=1$, as the ratio $L_{n,3,j}^{(u)}/\mathcal{D}_{1n}$ approaches 0, we have
@@
Similarly, the condition $L_{n,2,j}/\mathcal{D}_{1n}\to 0$ implies that $\dfrac{\sum_{i\in\mathcal{V}_{j}}\exp(b_{n,i})\|\Delta W_{n,2ij}\|_{1}}{\mathcal{D}_{1n}}\to 0$. Combining these results, we obtain